ework ? Saved Help Saw A family friend has asked your help in analyzing the operations of three anonymous companies operating in the same service sector industry. Supply the missing data in the table below: (Loss amounts should be indicated by a minus sign. Round your percentage answers to nearest whole percent) Company Sales Net operating income Average operating assets Return on investment (ROI) Minimum required rate of return s 370,000 740,000 S51,000 500,000 S 164,000 S 152.000 251% 191% 181% 12% Percentage Dollar amount S48,000 8,000 Residual income KPrey11 of 17 E Next >
Answer #1

Answer:

Company

A

B

C

Sales

370,000

740,000

500,000

Net operating income

41000

51000

26240

Average operating assets

164000

268421

152000

Return On ivestment (ROI)

25%

19%

17.26%

Minimum required return

Percentage

18%

17.88%

12%

Dollar amount

29520

48000

18240

Residual Income

11480

3000

8000

Working notes for the above answer is as under

Company A

1

Return on investment

=Net operating income/Average operating assets

25%= Net operating income/164000

Net operating income=164000*25%

Net operating income=41,000

2

Minimum required return in dollar

= Average operating assets* Minimum required return in %

=164,000*18%

=29520

3

Residual Income

=41000-29250

=11480

________________________________

Company B

Average operating assets

=51000/19%
=268421

Minimum required return in %

=Minimum required return in dollar/Average operating assets

=480000/268421

=17.88%

________________________________


465 466 468 469 470 471 472 473 474 475 476 477 Company Sales Net operating income Average operating assets Return On ivestme